Announcments & Equipment Logs

December, 2012

Facility back in action

The air supply has been restored, and the NMRs are ready for use.  

Posted in Uncategorized | No Comments »

Facility Shutdown

Due to the unannounced utility shutdown the NMR facility is now closed until further notice.  I have been told the air supply will be back up sometime between 11am and 3pm.  I will post when the air supply has returned and the facility is back in working order.

Posted in Uncategorized | No Comments »


The chiller unit is malfunctioning and therefore variable temperature is not available on the instrument until further updates.  All spectra need to be recorded at 298 K i.e. room temperature.

Posted in Hardware, Spectrometer | No Comments »

AV301 : PROBE UPDATE – 19F probe back

The erstwhile damaged 19F probe has been repaired in house and has been tested to function satisfactorily.  We thank the users for their patience and understanding while waiting for the 19F probe to be available.  Please let me or Paul know if you observe anything of significance with the TXD 1H {19F,13C} probe.

Posted in Hardware, Spectrometer | No Comments »